A company decides to establish an EOQ for an item. The annual demand is 400,000 units, each costing $8, ordering costs are $32 per order, and inventory-carrying costs are 20%. Calculate the following:

a. The EOQ in units

b. Number of orders per year

c. Cost of ordering

d. Cost of carrying inventory

e. Total Cost
Answer #1

Given values:

Annual demand (D) = 400,000 units

Cost per unit (C) = $8

Ordering costs (Co) = $32 per order

Inventory-carrying costs (Cc) = 20% of cost per unit = 20% of $8

Inventory-carrying costs (Cc) = $1.6

Solution:

(a) Economic Order Quantity (EOQ):

EOQ = SQRT [(2 x D x Co) / Cc]

EOQ = SQRT [(2 x 400,000 x $32) / $1.6]

EOQ = 4,000

Economic order quantity (EOQ) = 4,000 units

(b) Number of orders:

Number of orders = (Annual demand / EOQ)

Number of orders = (400,000 / 4,000)

Number of orders = 100

(c) Cost of ordering:

Annual ordering cost = (D / EOQ) x Co

Annual ordering cost = (400,000 / 4,000) x $32

Annual ordering cost = $3,200

(d) Cost of carrying inventory:

Annual inventory carrying cost = (EOQ / 2) x Cc

Annual inventory carrying cost = (4,000 / 2) x $1.6

Annual inventory carrying cost = $3,200

(e) Total cost:

Total cost = Annual ordering cost + Annual inventory carrying cost

Total cost = $3,200 + $3,200

Total cost = $6,400